I need help with a problem. The problem is 3qt.=_____ gal.
Found 3 solutions by nerdybill, Mathtut, solver91311:
Answer by nerdybill(7384)

Since there are 4 quarts per gallon
we can convert the 3 quarts into gallons by multiplication:
3 quarts * 1 gallon/4 quarts (notice the "quarts" cancel)
= 3/4 gallon

Answer by Mathtut(3670)

well since 4 quarts = 1 gallon
:
multiply both terms by 3/4 to get to 3 quarts
:
3/4(4 quarts)=3/4(1 gallon)
:
then 3 quarts = 3/4 gallon
:
another way to look at this is since 4 quarts makes 1 gallon then
:
1 quart makes 1/4gal
:
now if you multiply both the quarts and gal by 3
:
3(1quart)=3(1/4)gal
:
3quarts = 3/4 gal

Answer by solver91311(24713)

A gallon contains 4 quarts (hence the name). So if the whole is divided into 4 parts, and you have 3 of them...